A health food store mixes granola that costs them $4 per pound and raisins that cost them $2 per pound together to make 25 pounds of raisin granola. How many pounds of each do they need to produce a mixture that costs them $80?
Answer by ptaylor(2198)

Let x= amount of granola
Then 25-x=amount of raisins
$4x+$2(25-x)=$80 simplifying, we get:
$4x+$50-$2x=$80 or
$4x-$2x=$80-$50
$2x=$30
x=15 lbs of granola
25-x=25-15=10 lbs of raisins
ck
4(15)+2(10)=80
60+20=80
80=80
Hope this helps. Happy holidays--ptaylor